Calculate the standard heat of formation of aluminum sulfide given the following information?
2Al2S3(s) + 9O2(g) ->2Al2O3(s)+ 6SO2(g)
delta H of f = -3684.2kj
(The standard heat of formation : for SO2(g) -296.8kj/mol; Al2O3(s)= -1676kj/mol; and O2(g)= 0.0kj/mol.)
Answer #1

2Al2S3(s) + 9O2(g) ->2Al2O3(s)+ 6SO2(g)

delta Hof = 2* delta Hof (Al2O3) + 6*delta Hof (SO2) -2*delta Hof (Al2S3) - 9*delta Hof (O2)
-3684.2 = 2* (-1676) + 6* (-296.8) - 2*delta Hof (Al2S3) - 9*0
2*delta Hof (Al2S3) = 2* (-1676) + 6* (-296.8) + 3684.2
delta Hof (Al2S3) = -724.3 KJ/mol
Answer: -724.3 KJ/mol